OpenStudy (anonymous):

@Mathfematics Help!!! Peter has a total of 47 dimes & quarters in his pocket. The total value of the coins is $9.50. A. Write a system of equations. B. How many dimes & quarters does he have?

OpenStudy (anonymous):

Let d = number of dimes and q=number of quarters. A. {d + q = 47, 10 d + 25 q = 950} B. {d = 15, q = 32}\[(15\ .10)+(32\ .25)=\text{$\$$9}.50 \]